OpenStudy (anonymous):

Find a number whose product with 4 is the same as its sum with 12.

OpenStudy (mattfeury):

\[4 * x = 12 + x\]solve for x

OpenStudy (mattfeury):

\[4*x - x = 12\] x = 4

OpenStudy (anonymous):

4*x=x+12 solving for x, we get 4x-x=12 3x=12 x=12/3 X=4

OpenStudy (anonymous):

you can prove it, this way: 4(4)=4+12 16=16

OpenStudy (anonymous):

thanks